Question 6 xx = 1 True False
Answer #1

Given statement: X AND X' = 1

Irrespective of the value of X (High or Low)

X AND X' = 0

Therefore given statement:

X AND X' = 1 is FALSE

_______________________________________

Case 1:

If X = 0 (low), then X' = 1 (high)

Therefore, 1 AND 0 = 0(low)

Case 2:

If X = 1 (high), then X' = 0 (low)

Therefore, 0 AND 1 = 0 (low)

Therefore the statement is low in all the cases hence proved.

___________________________________________

ANSWER:

FALSE

______________________________________________

Feel free to ask any questions in the comments section
Thank You!